If my friend insist that there are more chances of getting a head on a single toss than a tail, my friend is wr
Need help with this assignment?Get an original answer from a qualified tutor — from $10/page.
Get it written →75 Word Response
Hi Class,
If my friend insist that there are more chances of getting a head on a single toss than a tail, my friend is wrong; the probability for any specific outcome is defined as a fraction or a proportion of all the possible outcomes (Gravetter, 2021). The probability of a specific outcome is expressed with a p (for probability), followed by the specific outcome in parentheses (Gravetter, 2021). When tossing a coin we use p(heads)=1/2=0.50=50%
All the possible probability values are contained in a limited range (Gravetter, 2021).
Get a plagiarism-free answer to this question
Send us your instructions and we’ll match you with the best writer in your subject.
- 100% human-written, zero AI
- Turnitin report included
- Confidential — we never share your data
- Free revisions & refunds